OpenStudy (daenio):

The moon's period of revolution around the Earth is 27.3 d. How far away is the moon? Question 5 options: A) 1.33× 107 m/s B) 3.83 × 108 m/s C) 3.77 × 108 m/s D) 4.88 × 109 m/s

OpenStudy (anonymous):

by applying kepler's law of planetary motion , you can find the answer. its easy

OpenStudy (daenio):

How do you do that?

OpenStudy (daenio):

@CarlosGP

OpenStudy (anonymous):

@daenio, something is wrong with the question as it is asking the distance however all the options are velocities. And I see 4 options, not 5 as announced

OpenStudy (anonymous):

Velocity = (2 pi R)/T T is the period and you are trying to find R

OpenStudy (anonymous):

You can plug that for the velocity into: F = G (Mm)/R^2 = mv^2/R

OpenStudy (anonymous):

Why are all of your answers in: m/s That is a velocity rather than a distance.

OpenStudy (daenio):

@CarlosGP Yeah, you're right. I totally forgot to change it. I notified my teacher earlier about the error.

OpenStudy (daenio):

@CarlosGP LOL, the question number is 5, not 5 options. Question 5.

OpenStudy (daenio):

Okay, I got B) for the answer!

OpenStudy (anonymous):

@daenio, have a look at this

OpenStudy (anonymous):

in VM should say RE and not RT

OpenStudy (daenio):

I think you're right. Not sure though. Anyways, as it turns out, C) is actually the answer.

OpenStudy (anonymous):

Answer would be B) if you consider distance from the center of the Earth. C) is the answer if you consider distance from the level of sea, that is B) minus Earth's ratius=3.83x10^8-0.0637x10^8=3.77x10^8

OpenStudy (anonymous):

Then B) or C) may be valid depending on where you are measuring from. I would take B) as it is usual to measure distance from center to center